# Towards Multi-agent Policy-based Directed Hypergraph Learning for Traffic Signal Control

### v1

@@ Abstract
Deep reinforcement learning (DRL) methods that incorporate graph neural networks (GNNs) have been extensively studied for intelligent traffic signal control, which aims to coordinate traffic signals effectively across multiple intersections. Despite this progress, the standard graph learning used in these methods still struggles to capture higher-order correlations in real-world traffic flow. In this paper, we propose a multi-agent proximal policy optimization framework DHG-PPO, which incorporates PPO and directed hypergraph module to extract the spatio-temporal attributes of the road networks. DHG-PPO enables multiple agents to ingeniously interact through the dynamical construction of hypergraph. The effectiveness of DHG-PPO is validated in terms of average travel time and throughput against state-of-the-art baselines through extensive experiments.
Index Terms:
Traffic signal control, directed hypergraph learning, multi-agent system, deep reinforcement learning.
I Introduction
Traffic Signal Control Systems (TSCs) are essential components of contemporary transportation infrastructure, playing a vital role in managing and coordinating the flow of traffic across road networks. Traditional traffic signal control methods rely on presetting light timings based on historical traffic patterns, periodically adjusted to manage vehicle flow through intersections[1]. However, the substantial rise in vehicle numbers has introduced challenges such as a lack of adaptability to dynamic traffic conditions and rudimentary delay coordination among traffic lights within the network. As a result, these systems fail to effectively mitigate traffic congestion throughout the road networks[2].
@@ I Introduction
On the other hand, recent advancements in Graph Neural Networks (GNNs) have improved the representation embeddings of traffic signal control nodes by utilizing the inherent topological structure of the road network [11]. Inspired by the idea of incorporating the reward feedback mechanism into the learning process, some recent studies have employed spatio-temporal graph neural network methods combined with multi-agent reinforcement learning to reduce vehicle travel time and alleviate congestion in the road network, yielding promising results[12, 13, 14]. However, these approaches do not comprehensively account for the impact of spatio-temporal traffic flow between adjacent intersections. Meanwhile, compressing complex relationships into pairwise relationships loses high-order spatial-temporal information, posing challenges in handling real-world dynamic road network situations. To address this issue, hypergraph learning offers a promising approach to capture the complex spatio-temporal relationships among neighboring intersections [15]. In contrast to the conventional graph in pairwise style, hypergraphs utilize hyperedges to connect more than two nodes to form hyperedges, by which provides distinct benefits in modeling complex relationships among multiple nodes and enable comprehensive higher-order network analysis [16, 17].
To this end, we propose a novel hypergraph-based multi-agent DRL framework for mining high-order hidden correlations among traffic lights in the road network. Moreover, we introduce Proximal Policy Optimization (PPO), which employs a simple yet effective clipping mechanism that restricts the magnitude of policy updates, preventing significant policy changes during the learning process. In large-scale complex environments, PPO is an effective solution among RL algorithms [18]. The main contributions of this paper are as below: (1) Framework: We propose a hypergraph-based multi-agent reinforcement learning framework. Agents are deployed at each intersection to coordinate traffic flow conditions between multiple intersections. Interaction between agents is realized by dynamically constructing directed hypergraphs. Combined with Multi-Agent Proximal Policy Optimization (MA-PPO), signal light decision optimization between multiple intersections strikes a balance between network-level and single-intersection-level. (2) Directed hypergraph learning: The directed hypergraph module is introduced to extract the directed spatial attributes of the road network according to the driving direction of traffic flow. Spatial-temporal correlations among multiple node groups are captured through hyperedges generated by a couple of dynamical headset and tailset, which contribute to updating node embeddings via the multi-head attention mechanism. (3) Experiments: We conduct rigorous experiments based on real-world road network datasets. The obtained results validate the effectiveness of our proposal compared with several state-of-the-art methods in terms of average travel time and throughput.
II Related Work
II-A Reinforcement Learning based methods for TSC
@@
Traditional TSC methods use deterministic offline strategies such as phase switching with predefined rules [19, 20, 21]. In contrast, reinforcement learning methods seek to adjust the signal light switching strategy adaptively based on the real-time situation at the intersection[22]. Although these DRL methods can create effective optimization strategies within individual intersections, the collaborative interaction between intersections is neglected. Therefore, when considering the overall road network, it may become trapped in local optimal situations. To mitigate this issue, max-pressure control [23, 24, 25], independent Advantage Actor-Critic (A2C) [26], and policy gradient method [27] are employed with multi-agent DRL schema, leading to a collaboration of various intersections for the global traffic optimization. However, the aforementioned studies primarily focus on extracting information from the first-order neighbors of each intersection, which overlooks the potential essential high-order information across different intersections.
II-B Integration of Reinforcement Learning and Graph Learning
@@
Due to the inherent topological properties of road networks, our proposed framework is closely related to the recent advances of GNN-based studies[13, 12, 28, 29]. Therein, Colight, proposed by Wei et al. [13], utilizes Graph Attention Networks (GAT) combined with parameter-sharing to balance the influences among adjacent intersections. Nishi et al. [12] demonstrate a model that uses Graph Convolutional Networks (GCN) to learn representations among multiple adjacent nodes. Although these two methods utilize the message-passing strategy in graph learning to absorb the representation information of adjacent nodes, the traffic flow direction on the graph are not incorporated. Comparatively, we present a directed hypergraph-based framework to enhance high-order correlation extraction among multiple nodes in the road networks.
III Methodology
III-A Problem definition
@@
This paper frames the control of traffic signal lights as a multi-agent reinforcement learning problem, deploying an intelligent agent at each intersection to manage the switching of signal light phases. To model the relationship between the traffic state and control action, it is reasonable to simplify the traffic signal control as the Markov decision process, and the effectiveness is supported by numerous empirical studies[5]. The problem is characterized by the following components:
@@
Action $\mathcal{A}$: We assume each intersection contains four signal phases: Phase 1 (going straight from West and East), Phase 2 (going straight from North and South), Phase 3 (turning left from West and East), Phase 4 (turning left from North and south). At time $t$, each agent $i$ selects one of four signal light phases as an action $a_{i}^{t}$.
@@
Reward $\mathcal{R}$ and Policy $\pi$ : After agent $i$ select the action, it obtains a reward $r_{i}^{t}$ by reward function $\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}\rightarrow\mathcal{R}$. We define the immediate reward for agent $i$ as $r_{i}^{t}=-\sum_{l}u_{i,l}^{t}$, where $u_{i,l}^{t}$ is the queue length on the approaching lane $l$ at time $t$. Agent $i$ chooses an action $a$ based on a certain policy $\mathcal{O}\times\mathcal{A}\rightarrow\pi$, aiming to maximize the total reward [13].
III-B Framework
We propose a novel framework based on MA-PPO, in which PPO is an on-policy gradient method for reinforcement learning [30]. Meanwhile, to effectively capture the spatio-temporal correlations among multiple sets of intersections, we introduce directed hypergraphs to enhance coordination among multiple agents deployed at intersections. In Figure 1, the module of directed hypergraph learning is consistent with the Spatial-Temporal Hypergraph Learning module proposed in [31].
Figure 1: An overview of our proposed framework.
III-B1 Directed Hypergraph
@@
Inspired by [31], we adopt the dynamic construction of directed hypergraphs to capture the spatio-temporal correlations between multiple sets of intersections by
@@
where $o_{i}^{t}\in\mathbb{R}^{d}$ is the observation of agent $i$ at time $t$. $W_{e}\in\mathbb{R}^{d\times l}$ and $b_{e}\in\mathbb{R}^{l}$ are trainable weight matrix and bias vector. $\sigma(\cdot)$ is the ReLU function. $h_{i}^{t}\in\mathbb{R}^{l}$ represents the traffic information of the $i$-th intersection at time $t$. ${H}_{\mathcal{G}}^{t}=\{h_{1}^{t},...,h_{N}^{t}\}\in R^{N\times l}$ is the initial node embedding of the entire road network.
Then, we capture the complex spatio-temporal correlations between nodes in the road network by dynamically constructing spatial hyperedges and temporal hyperedges. The spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$ can be generated based on the reconstruction of the master node $\dot{v}^{t}_{i}\in\emph{V}^{t}$ and the spatial candidate slave node set $\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)=\left\{v|v\in\emph{V}^{t},v\neq\dot{v}^{t}_{i}\right\}$, which is denoted as:
@@
where ${c}_{spa}\left(\dot{v}^{t}_{i}\right)$ denotes the spatial reconstruction error. $\left\|\cdot\right\|_{2}$ denotes the $l2$ norm of the vector. $\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)$ and $\emph{H}_{\mathcal{G}}^{t}\left(\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)\right)$ are node feature matrices of the master node and the spatial candidate slave node set respectively. $\theta_{spa}$ is a trainable projection matrix when generating the spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$. $\emph{p}^{spa}_{\dot{v}^{t}_{i}}\in\mathcal{R}^{\left(N-1\right)}$ is the trainable reconstruction coefficient vector. $S^{spa}\left(\dot{v}^{t}_{i}\right)$ is denoted as:
$S^{spa}\left(\dot{v}^{t}_{i}\right)=\left\{v|v\in\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right),\emph{p}^{spa}_{\dot{v}^{t}_{i}}\left(v\right)>0\right\}$ (3)
Based on the trainable vector $\emph{p}^{spa}_{\dot{v}^{t}_{i}}$, we filter the slave nodes to construct the tail node set $\mathcal{H}_{T}$ in a more flexible manner. We group master nodes that share the same slave node set to form the head node set $\mathcal{H}_{H}$. Finally, we form directed hyperedges composed of head and tail node-sets. Similarly, temporal slave nodes are encircled from $\emph{V}^{t-1}$ to form the temporal hyperedge. Overall, the loss of hyperedges generation in one timestamp is defined as:
@@
where $\left\|\cdot\right\|_{1}$ denotes the $l1$ norm of the vector and $c_{tem}$ denotes the reconstruction error of temporal hyperedges. $\lambda$ is the weight hyperparameter. $\gamma_{2}$ is the regularizing factor to balance $l1$ norm and $l2$ norm.

### v2

# DHLight: Multi-agent Policy-based Directed Hypergraph Learning for Traffic Signal Control

@@ Abstract
Recent advancements in Deep Reinforcement Learning (DRL) and Graph Neural Networks (GNNs) have demonstrated notable promise in the realm of intelligent traffic signal control, facilitating the coordination across multiple intersections. However, the traditional methods rely on standard graph structures often fail to capture the intricate higher-order spatio-temporal correlations inherent in real-world traffic dynamics. Standard graphs cannot fully represent the spatial relationships within road networks, which limits the effectiveness of graph-based approaches. In contrast, directed hypergraphs provide more accurate representation of spatial information to model complex directed relationships among multiple nodes. In this paper, we propose DHLight, a novel multi-agent policy-based framework that synergistically integrates directed hypergraph learning module. This framework introduces a novel dynamic directed hypergraph construction mechanism, which captures complex and evolving spatio-temporal relationships among intersections in road networks. By leveraging the directed hypergraph relational structure, DHLight empowers agents to achieve adaptive decision-making in traffic signal control. The effectiveness of DHLight is validated against state-of-the-art baselines through extensive experiments in various network datasets. We release the code to support the reproducibility of this work at https://github.com/LuckyVoasem/Traffic-Light-control
1 Introduction
Traffic Signal Control Systems (TSCSs) are essential components of contemporary transportation infrastructure, playing a vital role in managing and coordinating the flow of traffic across road networks. Traditional traffic signal control methods rely on presetting light timings based on historical traffic patterns, periodically adjusted to manage vehicle flow through intersections [7]. However, the substantial rise in vehicle numbers has introduced challenges such as a lack of adaptability to dynamic traffic conditions and rudimentary delay coordination among traffic lights within the network. As a result, these systems fail to effectively mitigate traffic congestion throughout the road networks [42, 17].
@@ 1 Introduction
On the other hand, recent advancements in Graph Neural Networks (GNNs) have improved the representation embeddings of traffic signal control nodes by utilizing the inherent topological structure of the road network [18]. Inspired by the idea of incorporating the reward feedback mechanism into the learning process, some recent studies have employed spatio-temporal graph neural network methods combined with multi-agent reinforcement learning to reduce vehicle travel time and alleviate congestion in the road network, yielding promising results [16, 31, 41]. However, these approaches do not comprehensively account for the impact of spatio-temporal traffic flow between adjacent intersections. Meanwhile, compressing complex relationships into pairwise relationships loses high-order spatial-temporal information, posing challenges in handling real-world dynamic road network situations. To address this issue, hypergraph learning offers a promising approach to capture the complex spatio-temporal relationships among neighboring intersections [1]. In contrast to the standard graph in pairwise style, hypergraphs utilize hyperedges to connect more than two nodes to form hyperedges, by which provides distinct benefits in modeling complex relationships among multiple nodes and enable comprehensive higher-order network analysis [14, 4]. However, hypergraphs may overlook the directed nature of spatio-temporal relationships, which limits their effectiveness in accurately capturing dynamic interactions in real-world scenarios like TSCSs.
To this end, we propose a novel multi-agent policy-based directed hypergraph learning DHLight for mining high-order hidden correlations among traffic lights in the road network. Compared to the traditional undirected hypergraphs, directed hypergraphs are superior in modeling the directionality of traffic flow and have a distinct advantage in representing complex directed relationships among multiple nodes exist in TSCSs. Moreover, we utilize the Multi-agent Proximal Policy Optimization (MA-PPOs [20]), which employs a simple yet effective clipping mechanism that restricts the magnitude of policy updates, preventing significant policy changes during the learning process. In large-scale complex environments, MA-PPO is an effective solution among DRL algorithms [12]. The main contributions of this paper are as below:
  • Framework: We propose a multi-agent policy-based directed hypergraph learning framework. In this framework, agents are deployed at each intersection to coordinate traffic flow conditions between multiple intersections. Interaction between agents is realized by dynamically constructing directed hypergraphs. Empowered by with MA-PPO, signal light decision optimization between multiple intersections strikes a balance between network-level and single-intersection-level.
  • Directed hypergraph learning: The directed hypergraph module is introduced to extract the directed spatial attributes of the road network according to the driving direction of traffic flow. Spatial-temporal correlations among multiple node groups are captured through hyperedges generated by a couple of dynamical headset and tailset, which contribute to updating node embeddings via the multi-head attention mechanism.
  • Experiments: We conduct rigorous experiments based on two real-world road network datasets and two synthetic network datasets. The obtained results validate the effectiveness of our proposal against several state-of-the-art methods in terms of average travel time.
The rest of this paper is organized as follows: Section 2 summarizes the related work. Section 3 formulates the traffic signal control as a multi-agent reinforcement learning problem, following by the description of our proposed multi-agent policy-based directed hypergraph learning framework in Section 4. Section 5 evaluates the performance of our proposed DHLight. Finally, the paper is concluded with future work in Section 6.
2 Related Work
2.1 Reinforcement Learning for TSCSs
@@
Traditional TSC methods use deterministic offline strategies such as phase switching with predefined rules [11, 10, 23]. In contrast, reinforcement learning methods seek to adjust the signal light switching strategy adaptively based on the real-time situation at the intersection [17]. Although these DRL methods can develop effective optimization strategies for individual intersections, the in-depth collaborative interaction between intersections remains inadequately explored. Therefore, when considering the overall road network, it may become trapped in local optimal situations. To mitigate this issue, max-pressure control [2, 30, 5], independent Advantage Actor-Critic (A2C) [6], policy gradient method [32], MixLight [33], and GPLight [13] are employed with multi-agent DRL schema, leading to a collaboration of various intersections for the global traffic optimization. However, the aforementioned studies primarily focus on extracting information from the first-order neighbors of each intersection, which overlooks the potential essential high-order information across different intersections.
2.2 Integration of Reinforcement Learning and Graph Learning
@@
Due to the inherent topological properties of road networks, our proposed framework is closely related to the recent advances of GNN-based studies [31, 16, 35, 40, 27, 24]. Therein, Colight [31] utilizes Graph Attention Networks (GAT) combined with parameter-sharing to balance the influences among adjacent intersections. Nishi et al. [16] demonstrated a model that uses Graph Convolutional Networks (GCN) to learn representations among multiple adjacent nodes. GPLight [13] aggregates neighborhood features from distant intersections using GCN and a grouping model. eMARLIN+ [27] achieves the sharing of embedded information among adjacent intersections through an encoder-decoder structure. Wang et al. [24] proposed a spatio-temporal hypergraph-based multi-agent reinforcement learning framework, which integrates hypergraph learning [38] into the multi-agent soft actor-critic algorithm to optimize traffic signal control. Although these methods utilize the message-passing strategy in graph learning to absorb the representation information of adjacent nodes, the traffic flow direction on the graph are not incorporated. Comparatively, we present a directed hypergraph-based framework to enhance high-order correlation extraction among multiple nodes in the road networks.
3 Problem Formulation
In this paper, we assume a road traffic scenario featuring three vehicle lanes and four directions of movement: East (E), South (S), West (W), and North (N). A traffic light is deployed at each intersection to regulate traffic flow by managing three types of movements: Drive Through (T), Turn Left (L), and Turn Right (R), achieved through a sequence of different phases. A phase consists of a set of vehicular movement signals that do not conflict with one another. Figure 1 illustrates a road intersection with four distinct phases: Phase 1 (ET and WT), Phase 2 (EL and WL), Phase 3 (ST and NT), and Phase 4 (NL and SL). Vehicles making a right turn at this intersection can proceed during any of these four phases. An agent is employed to manage the phase changes at each intersection. At any given time $t$, the agent assesses the current intersection conditions and chooses one of the four available phases.
@@ 3 Problem Formulation
This paper frames the control of traffic signal lights as a multi-agent reinforcement learning problem, which deploys an intelligent agent at each intersection to manage the switching of signal light phases. To model the relationship between the traffic state and control action, it is reasonable to simplify the traffic signal control as the Markov decision process, and the effectiveness is supported by numerous empirical studies[29]. The problem is characterized by the following components:
@@
Action $\mathcal{A}$: We assume each intersection contains assumed four signal phases. At each time $t$, each agent $i$ chooses one of the four traffic signal phases as its action $a_{i}^{t}$. The decisions made by the traffic signal at an intersection can impact the traffic flow in the surrounding area, which subsequently affects the observations of the agent.
@@
Reward $\mathcal{R}$ and Policy $\pi$ : After agent $i$ select the action, it obtains a reward $r_{i}^{t}$ by reward function $\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{N}\rightarrow\mathcal{R}$. We define the immediate reward for agent $i$ as:
$r_{i}^{t}=\text{-}\sum_{l}{u_{i,l}^{t}}$ (1)
where $u_{i,l}^{t}$ is the queue length on the approaching lane $l$ at time $t$.At each intersection, the waiting time for vehicles is affected by the queue length in the lane. As the queue becomes longer, the waiting time for vehicles increases, leading to an overall increase in travel time. Agent $i$ chooses an action $a$ based on a certain policy $\mathcal{O}\times\mathcal{A}\rightarrow\pi$, with the goal of reducing travel time for all vehicles within the road network.
Figure 1: An overview of our proposed framework DHLight.
4 Multi-agent Policy-based Directed Hypergraph Learning Framework
In this section, we introduce our proposed framework DHLight for TSCSs. This section first gives an overview of DHLight. Then, a detailed description of the two key modules is presented: directed hypergraph module (including hyperedge construction and directed hyperedge learning) and multi-agent deep reinforcement learning module. The time complexity analysis is provided at the end of this section.
4.1 Overview
We propose a novel framework DHLight to extract the directed spatio-temporal attributes of the road networks. As illustrated in Figure 1, the overview of our proposed DHLight is summarized as below:
Firstly, we obtain real-time intersection status from the agents deployed at each intersection as the input to the network, which refers to the number of vehicles in each lane of the intersection and the encoding of the current phase of the traffic lights. Subsequently, we employ an on-policy gradient reinforcement learning method PPO, to continuously optimize decision-making at each intersection, thereby enhancing policy performance, as depicted in the bottom right corner of Figure 1. The PPO composes of an actor network and a critic network. In the actor network, by using the intersection’s status, we calculate the probability distribution of each action, then stochastically select an action based on this distribution for execution, thereby enabling the phase transition at the current intersection. Meanwhile, in the critic network, to effectively capture spatio-temporal relationships across different intersections, we utilize the directed hypergraphs to enhance coordination among multiple agents at these intersections. As illustrated in the top right section of Figure 1, we introduce a directed hypergraph module (comprising directed hyperedge construction and directed hyperedge learning), aimed at capturing the complex and directed spatio-temporal relationships among multiple nodes in traffic signal control, thus updating the representation of each node. Finally, the critic network optimizes the actor network process to enhance decision-making efficiency, thereby improving the overall policy performance of TSCSs. Moreover, when dealing with large-scale road networks, a partitioning approach can be adopted to divide them into multiple sub-networks, with each sub-network being controlled by DHLight respectively.
4.2 Directed Hypergraph Construction
We assume that the spatial-temporal hypergraph at time $t$ is represented as $\mathcal{G}^{t}=\left\{\mathcal{V}^{t},\mathcal{E}^{t}\right\}$. To account for the temporal relationships among nodes, we incorporate the features of nodes from the previous timestamp to capture historical data following the method proposed in [37]. Consequently, the node set $\mathcal{V}^{t}$ includes nodes from the two consecutive timestamps, defined as $\mathcal{V}^{t}=\left\{\emph{V}^{t},\emph{V}^{t-1}\right\}$. The feature matrices for these timestamps are denoted as $\emph{X}^{t}$ and $\emph{X}^{t-1}$, respectively.
A hyperedge encompasses a specific group of nodes that share common attributes or implicit data relationships, representing local group dynamics among the nodes. To facilitate the creation of hyperedges, we introduce two types of nodes: master nodes and candidate nodes. A master node $\dot{v}\in\mathcal{V}$ serves as a reference point when forming the hyperedge $e\left(\dot{v}\right)\in\mathcal{E}$, which is formed in conjunction with a collection of candidate nodes $S\left(\dot{v}\right)=\left\{\hat{v}\right\}$. We propose a dynamic learning framework that generates two categories of hyperedges: the spatial hyperedges, which capture heterogeneity by encoding relationships among different channels at a single timestamp, and the temporal hyperedges, which examine interactivity by modeling the ongoing interactions of channels across consecutive timestamps.
@@ 4.2 Directed Hypergraph Construction
Inspired by the methodology proposed in [9] designed for hypergraph construction, we adopt the dynamic construction of directed hypergraphs to capture the spatio-temporal correlations between multiple sets of intersections by:
@@
where $o_{i}^{t}\in\mathbb{R}^{d}$ is the observation of agent $i$ at time $t$, which incorporates the traffic flow state of the intersection and the traffic signal phase. $W_{e}\in\mathbb{R}^{d\times l}$ and $b_{e}\in\mathbb{R}^{l}$ are trainable weight matrix and bias vector. $\sigma(\cdot)$ is the ReLU function. $h_{i}^{t}\in\mathbb{R}^{l}$ represents the traffic information of the $i$-th intersection at time $t$. ${H}_{\mathcal{G}}^{t}=\{h_{1}^{t},...,h_{N}^{t}\}\in R^{N\times l}$ is the initial node embedding of the entire road network.
Then, we capture the complex spatio-temporal correlations between nodes in the road network by dynamically constructing spatial hyperedges and temporal hyperedges. The spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$ can be generated based on the reconstruction of the master node $\dot{v}^{t}_{i}\in\emph{V}^{t}$ and the spatial candidate node set $\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)=\left\{v|v\in\emph{V}^{t},v\neq\dot{v}^{t}_{i}\right\}$, which is denoted as:
@@
where ${c}_{spa}\left(\dot{v}^{t}_{i}\right)$ denotes the spatial reconstruction error. $\left\|\cdot\right\|_{2}$ denotes the $l2$ norm of the vector. $\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)$ and $\emph{H}_{\mathcal{G}}^{t}\left(\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right)\right)$ are node feature matrices of the master node and the spatial candidate node set respectively. $\theta_{spa}$ is a trainable projection matrix when generating the spatial hyperedge $e_{spa}\left(\dot{v}_{i}^{t}\right)\in\mathcal{E}^{t}$. $\emph{p}^{spa}_{\dot{v}^{t}_{i}}\in\mathcal{R}^{\left(N-1\right)}$ is the trainable reconstruction coefficient vector. $S^{spa}\left(\dot{v}^{t}_{i}\right)$ is denoted as:
$S^{spa}\left(\dot{v}^{t}_{i}\right)=\left\{v|v\in\tilde{S}^{spa}\left(\dot{v}^{t}_{i}\right),\emph{p}^{spa}_{\dot{v}^{t}_{i}}\left(v\right)>\zeta\right\}$ (4)
Based on the trainable vector $\emph{p}^{spa}_{\dot{v}^{t}_{i}}$ and threshold $\zeta$, we filter the candidate nodes to construct the tail node set $\mathcal{H}_{T}$ in a more flexible manner. We group master nodes that share the same candidate node set to form the head node set $\mathcal{H}_{H}$. Finally, we form directed hyperedges composed of head and tail node-sets. Similarly, temporal candidate nodes are encircled from $\emph{V}^{t-1}$ to form the temporal hyperedge. Overall, the loss of hyperedges generation in one timestamp is defined as:
@@
where $\left\|\cdot\right\|_{1}$ denotes the $l1$ norm of the vector and $c_{tem}$ denotes the reconstruction error of temporal hyperedges. $\lambda$ is the weight hyperparameter. $\gamma_{2}$ is the regularizing factor to balance $l1$ norm and $l2$ norm. The reconstruction error ensures accurate feature restoration, $l1$ aids feature selection, and $l2$ prevents overfitting, enabling the model to capture high-order spatio-temporal correlations in road networks.
4.3 Directed Hypergraph Learning